1) The Cobb-Douglas production function is 

y=z1αz2β


, with 

α>0


 and 

β>0


. Show that 

MP1=αy/z1


MP2=βy/z2


. What is the 

MRTS21


? How does it vary with: (a) y; (b) 

z2/z1


.


1
Expert's answer
2022-03-24T16:11:14-0400

"Y=Z_1^{\alpha}Z_2^{\beta}\\
MP_1= \alpha Z_1^{\alpha -1}Z_2^{\beta}\\
MP_1= \frac{\alpha Z_1^{\alpha }Z_2^{\beta}}{Z_1}\\
MP_1= \frac{\alpha Y}{Z_1}\\"


"MP_2= \beta Z_1^{\alpha }Z_2^{\beta-1}\\
MP_2= \frac{\beta Z_1^{\alpha }Z_2^{\beta}}{Z_2}\\
MP_2= \frac{\beta Y}{Z_2}\\"


"MRTS= \frac{MP_1}{MP_2}\\
MRTS= \frac{\frac{\alpha Y}{Z_1}}{\frac{\beta Y}{Z_2}}\\
MRTS= \frac{\alpha}{\beta}*\frac{Z_2}{Z_1}\\"


a. A change in Y would not affect MRTS since MRTS is not a function of Y.


b. As "\frac{Z_2}{Z_1}" changes MRTS changes.

An increase in "\frac{Z_2}{Z_1}" MRTS increases and vice versa